OpenStudy (anonymous):

-5,5-5,5… A. subtract 10 from the previous term; -5,5 B. multiply the previous term by 1; 5 , 5 C. add 10 to the previous term; -5,-15 D. multiply the previous term by -1;-5, 5

OpenStudy (anonymous):

@iGreen. @adilalvi @Love_Ranaa @I_Always_Smiling @Asadkarim7 @blueSurrency

OpenStudy (cwrw238):

Hint; its a multiplication how do you get from -5 to 5 then from 5 to -5 ?

OpenStudy (anonymous):

Yep, divide them. -5 / 5 = ? 5 / -5 = ?

OpenStudy (anonymous):

Can you divide those? @appleandgraphphone

OpenStudy (anonymous):

who it be -0 >?

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

If it was 0..all the values would be 0

OpenStudy (anonymous):

-5 , 5 ? @iGreen.

OpenStudy (anonymous):

-5 / 5 = ? 5 / -5 = ? @appleandgraphphone Divide these and you get your answer.

OpenStudy (anonymous):

Yes, that is the result

OpenStudy (anonymous):

So the answer is either A or B

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

Hmm..can you only pick 1 answer?

OpenStudy (anonymous):

only 1

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

Hmm..I believe A is incorrect though.

OpenStudy (anonymous):

Oh yeah, nevermind.. -5 - 10 = -15, it should equal 5, but it doesn't. So A is incorrect.

OpenStudy (anonymous):

Understand?

OpenStudy (anonymous):

wait so its B > @iGreen.

OpenStudy (anonymous):

No, we already eliminated B and C.

OpenStudy (anonymous):

i meant D sorry .

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

Yep, D is correct.

OpenStudy (anonymous):

-5 * -1 = 5 5 * -1 = -5 -5 * -1 = 5 5 * -1 = -5

OpenStudy (anonymous):

\(\bbox [10pt, lime, border:5pt solid black]{\Huge\cal\color{red}\diamondsuit\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\diamondsuit}}}\)